OpenStudy (anonymous):

I need a genius. This a a university entrance exam's question: Find x^-1 . This is algebra. A graphing calculator is not allowed.

OpenStudy (nubeer):

hmm is that all the information given/

OpenStudy (anonymous):

Here's the equation \[x^{x^{\frac{1}{2}}} = \frac{1}{\sqrt{2}}\] I need to find: \[x^{-1}\]

OpenStudy (anonymous):

remember the exponential rule: \[a^{-m}={1\over a^m}\] therefore, \[x^{-1}={1\over x}\] for each number, take the reciprocal

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@terenzreignz

OpenStudy (amistre64):

\[x^{x^{\frac{1}{2}}} = \frac{1}{\sqrt{2}}\] \[log(x^{x^{\frac{1}{2}}}) = log(\frac{1}{\sqrt{2}})\] \[x^{1/2}log(x) = log(1)-log(\sqrt{2})\] \[x^{1/2}log(x) = -log(\sqrt{2})\] hmmm, not sure if this is gonna pan out this way

OpenStudy (anonymous):

To be honest I don't think log is needed since it's a question from Chapter 2. It's all about exponents.

OpenStudy (amistre64):

x^(1/2) = -1 might be what they are asking for then

OpenStudy (amistre64):

is that\[x^{(x^{1/2})}~or~(x^x)^{1/2}\]

OpenStudy (anonymous):

Without parenthesis

terenzreignz (terenzreignz):

They both have parentheses?

OpenStudy (amistre64):

they are not the same .... http://www.wolframalpha.com/input/?i=y%3D%28x%5Esqrt%28x%29%2C+y+%3D+sqrt%28x%5Ex%29

OpenStudy (anonymous):

the first one.

OpenStudy (anonymous):

the second one is totally different

OpenStudy (amistre64):

then for the first one, id assume that x^(-1) = x^(sqrt(x)), when -1 = sqrt(x)

OpenStudy (amistre64):

but thats not taking into account the ... = sqrt(2)/2 stuff

OpenStudy (anonymous):

\[x^{x^{1/2}}={1\over\sqrt{2}}\\ \implies \sqrt{2}\times x^{x^{1/2}}=1\\ \implies x^{1/2}=0\\ \implies x=0 \] the given equation cannot be satisfied

OpenStudy (anonymous):

atleast not for \(x\in\mathcal{R}\)

terenzreignz (terenzreignz):

How exactly does this follow? \[x^{x^{1/2}}={1\over\sqrt{2}}\\ \implies \sqrt{2}\times x^{x^{1/2}}=1\\ \color{red}{\implies x^{1/2}=0}\\ \implies x=0\]

OpenStudy (amistre64):